OpenStudy (anonymous):

Together, Steve and Tom sold 101 raffle tickets for their school. Steve sold 17 more than three times as many raffle tickets as Tom. How many did each boy sell?

OpenStudy (anonymous):

s + t = 101 s = 17+3t do you understand how I got the equations?

OpenStudy (anonymous):

for ease - substitute the value of s (2nd equation) into the 1st equation...solve for t

OpenStudy (anonymous):

Yeah, I used substitution I think I just made an algebraic error! Thanks again!

OpenStudy (anonymous):

17 + 3t + t = 101 4t = 101-17 4t = 84 t = 21

OpenStudy (anonymous):

now since we have T I will use the 2nd equation to determine S s = 17 + 3(21) s = 17 + 63 s = 80 now check : s+ t = 101 80 + 21 = 101

OpenStudy (anonymous):

Thanks! :D